You own a high-tech manufacturing entity. You would like to expand your operations but to do so you need to either lease or buy a $1.2 million piece of equipment for the next three years. The lease payments would be $475,000 a year for the three years. If the equipment is purchased, it will be depreciated straight-line to zero over the three-year period. The equipment will have no residual value at the end of the three years. Should the equipment be leased, the lessor and the lessee will both have marginal tax rates of 34%. The loan rate for your firm for this purpose is 8% pre-tax.

What is the amount of the annual after-tax lease payment to the lessee?

Multiple Choice

  • $449,920

  • $313,500

  • $475,000

  • $104,500

  • $209,600
Answer #1

Lessee can have tax benefit of lease payment, so for lessee after tax lease payment will be,

ANNUAL AFTER TAX LEASE PAYMENT = LEASE PAYMENT*(1-t) = 475000*(1-0.34) = 313500

Answer : $ 313500 (Thumbs up please)